# Ravan: Multi-Head Low-Rank Adaptation for Federated Fine-Tuning

### v1

@@ A.3 Experimental Settings
ViT-B-16/CIFAR-100 ViT-B-16/SVHN T5-Base/20 Newsgroups T5-Base/MRQA LLaMA3.2-1B/GLUE
@@
Communication Rounds $50$ $50$ $100$ $20$ $20$
Total Epochs (per round) $1$ $1$ $1$ $1$ $1$
[…]
@@
Method ViT-B-16/CIFAR-100 ViT-B-16/SVHN T5-Base/20 Newsgroups T5-Base/MRQA LLaMA3.2-1B/GLUE
FedIT $5{\times}10^{-3}$ $1{\times}10^{-3}$ $1{\times}10^{-3}$ $1{\times}10^{-3}$ $1{\times}10^{-4}$
FedEx-LoRA $1{\times}10^{-3}$ $1{\times}10^{-3}$ $1{\times}10^{-3}$ $1{\times}10^{-3}$ $1{\times}10^{-4}$
FFA-LoRA $1{\times}10^{-2}$ $1{\times}10^{-2}$ $1{\times}10^{-2}$ $5{\times}10^{-3}$ $1{\times}10^{-3}$
Fed-SB $1{\times}10^{-3}$ $5{\times}10^{-3}$ $5{\times}10^{-3}$ $1{\times}10^{-3}$ $1{\times}10^{-4}$
@@
Method ViT-B-16/CIFAR-100 ViT-B-16/SVHN T5-Base/20 Newsgroups T5-Base/MRQA LLaMA3.2-1B/GLUE
FedIT $5{\times}10^{-3}$ $1{\times}10^{-3}$ $1{\times}10^{-3}$ $5{\times}10^{-4}$ $1{\times}10^{-4}$
FedEx-LoRA $1{\times}10^{-3}$ $1{\times}10^{-3}$ $1{\times}10^{-3}$ $5{\times}10^{-4}$ $1{\times}10^{-4}$
FFA-LoRA $1{\times}10^{-2}$ $1{\times}10^{-2}$ $1{\times}10^{-2}$ $1{\times}10^{-2}$ $1{\times}10^{-3}$
Fed-SB $1{\times}10^{-3}$ $1{\times}10^{-3}$ $5{\times}10^{-3}$ $5{\times}10^{-4}$ $1{\times}10^{-4}$
RAVAN $5{\times}10^{-4}$ $5{\times}10^{-4}$ $1{\times}10^{-4}$ $1{\times}10^{-4}$ $5{\times}10^{-5}$
Table 10: Optimal learning rates – lower parameter budget / non-I.I.D. clients.
@@
Method ViT-B-16/CIFAR-100 ViT-B-16/SVHN T5-Base/20 Newsgroups T5-Base/MRQA
FedIT $5{\times}10^{-3}$ $1{\times}10^{-3}$ $1{\times}10^{-3}$ $1{\times}10^{-3}$
FedEx-LoRA $1{\times}10^{-3}$ $1{\times}10^{-3}$ $1{\times}10^{-3}$ $1{\times}10^{-3}$
FFA-LoRA $1{\times}10^{-2}$ $1{\times}10^{-2}$ $1{\times}10^{-2}$ $5{\times}10^{-3}$
Fed-SB $5{\times}10^{-4}$ $5{\times}10^{-3}$ $1{\times}10^{-3}$ $5{\times}10^{-4}$
@@
Table 11: Optimal learning rates – higher parameter budget / non-I.I.D. clients.
@@
Method ViT-B-16/CIFAR-100 ViT-B-16/SVHN T5-Base/20 Newsgroups T5-Base/MRQA
FedIT $5{\times}10^{-3}$ $1{\times}10^{-3}$ $1{\times}10^{-3}$ $5{\times}10^{-4}$
FedEx-LoRA $1{\times}10^{-3}$ $5{\times}10^{-4}$ $1{\times}10^{-3}$ $5{\times}10^{-4}$
FFA-LoRA $1{\times}10^{-2}$ $1{\times}10^{-2}$ $1{\times}10^{-2}$ $5{\times}10^{-3}$
Fed-SB $5{\times}10^{-4}$ $1{\times}10^{-3}$ $1{\times}10^{-3}$ $5{\times}10^{-4}$
RAVAN $5{\times}10^{-4}$ $5{\times}10^{-4}$ $1{\times}10^{-4}$ $1{\times}10^{-4}$
@@
Figure 5: Fraction of clients assigned to each trainable parameter budget in each distribution. Skewed-left is omitted because it is never used in our experiments.
To emulate computational heterogeneity in our FL setup, we vary the number of trainable parameters at each client. Let $N_{\max}$ denote the largest number of trainable parameters that any client can afford. Each client $c$ is constrained to a trainable parameter budget $N_{c}\in\{\frac{1}{4},\frac{1}{2},\frac{3}{4},1\}N_{\max}$. This value is held constant throughout the FL procedure to mirror fixed hardware limits. $N_{c}$ simply determines the trainable parameter budget per-client; all other hyperparameters are identical to the compute-homogeneous experiments. The bar plot in Figure 5 provides a visual summary of the client mix. Uniform serves as a neutral baseline where there is an equal proportion of clients at every trainable parameter budget; bell-shaped concentrates clients around medium ranks, reflecting the case where most devices have moderate capability; skewed-right stresses the system by placing a large share of the population at the lowest rank, leaving only a small fraction of high-capacity contributors. HetLoRA, FlexLoRA, and Ravan each accommodate clients with different trainable-parameter budgets in distinct ways:
  • HetLoRA and FlexLoRA: The LoRA rank for each client $c$ is $r_{c}=\bigl{\lfloor}\left(N_{c}\,/\,N_{\max}\right)\cdot r_{\max}\bigr{\rfloor}$ where $r_{\max}$ is the maximum rank trained by any client. Since the number of trainable parameters scales linearly with the rank, this scaling ensures that every client keeps its update within the allotted budget $N_{c}$ while allowing higher-capacity devices to contribute proportionally higher-rank updates.
@@
ViT-B-16 query,  value
T5-Base SelfAttention.q,  SelfAttention.v
LLaMA3.2-1B q_proj,  v_proj
For every model backbone, we insert LoRA adapters only in the self-attention projection matrices. The exact parameters for which we apply LoRA are described in Table 12. All other parameters are frozen and do not have associated LoRA parameters.
@@ A.4 Additional Experiments
(a) CIFAR-100 / lower parameter budget
(b) CIFAR-100 / higher parameter budget
@@
Figure 6: Impact of computational heterogeneity on baselines and Ravan across four datasets. Each row groups two-budget plots for the datasets listed in order.
@@
Method MNLI-MM MNLI-M QNLI QQP SST-2 RTE Average
@@
Method MNLI-MM MNLI-M QNLI QQP SST-2 RTE Average
@@
Method MNLI-MM MNLI-M QNLI QQP SST-2 RTE Average
@@
Method MNLI-MM MNLI-M QNLI QQP SST-2 RTE Average
@@
(a) CIFAR-100 / 20 clients / lower parameter budget
(b) CIFAR-100 / 50 clients / lower parameter budget
(c) CIFAR-100 / 20 clients / higher parameter budget
(d) CIFAR-100 / 50 clients / higher parameter budget
@@
Figure 7: Federated-learning training curves for CIFAR-100 and SVHN under different client counts and parameter-budget settings.

### v2

@@ A.3  Experimental Settings
ViT-B-16/CIFAR-100 ViT-B-16/SVHN T5‑Base/20 Newsgroups T5‑Base/MRQA LLaMA3.2‑1B/GLUE
@@
Communication Rounds $50$ $50$ $100$ 2020 2020
Total Epochs (per round) 11 11 11 11 11
For Ravan and each baseline, we run a learning rate hyperparameter sweep across the values $\{5\mathrm{e}{-5},\ 1\mathrm{e}{-5},\$
$5\mathrm{e}{-4},\ 1\mathrm{e}{-4},\ 5\mathrm{e}{-3},\ 1\mathrm{e}{-3},\ 5\mathrm{e}{-2},\ 1\mathrm{e}{-2},\ 5\mathrm{e}{-2}\}$ and choose the most performant learning to represent in our results. The following tables represents the optimal choices for each baseline in all settings. The following results each use the Adam optimizer with momentum set to $0.9$.
@@
Method ViT‑B‑16/CIFAR‑100 ViT‑B‑16/SVHN T5‑Base/20 Newsgroups T5‑Base/MRQA LLaMA3.2‑1B/GLUE
FedIT $5{\times}10^{-3}$ ${\times}10^{-3} 1×1$${\times}10^{-3}$ ${\times}10^{-3} 1×1$${\times}10^{-4}$
FedEx‑LoRA ${\times}10^{-3} 1×1$${\times}10^{-3}$ ${\times}10^{-3} 1×1$${\times}10^{-3}$ $1{\times}10^{-4}$
FFA‑LoRA ${\times}10^{-2} 1×1$${\times}10^{-2}$ $1{\times}10^{-2}$ $5{\times}10^{-3}$ $1{\times}10^{-3}$
Fed‑SB $1{\times}10^{-3}$ $5{\times}10^{-3}$ $5{\times}10^{-3}$ ${\times}10^{-3} 1×1$${\times}10^{-4}$
@@
Method ViT‑B‑16/CIFAR‑100 ViT‑B‑16/SVHN T5‑Base/20 Newsgroups T5‑Base/MRQA LLaMA3.2‑1B/GLUE
FedIT $5{\times}10^{-3}$ ${\times}10^{-3} 1×1$${\times}10^{-3}$ $5{\times}10^{-4}$ $1{\times}10^{-4}$
FedEx‑LoRA ${\times}10^{-3} 1×1$${\times}10^{-3}$ $1{\times}10^{-3}$ $5{\times}10^{-4}$ $1{\times}10^{-4}$
FFA‑LoRA ${\times}10^{-2} 1×1$${\times}10^{-2}$ ${\times}10^{-2} 1×1$${\times}10^{-2}$ $1{\times}10^{-3}$
Fed‑SB ${\times}10^{-3} 1×1$${\times}10^{-3}$ $5{\times}10^{-3}$ $5{\times}10^{-4}$ $1{\times}10^{-4}$
RAVAN $5{\times}10^{-4}$ $5{\times}10^{-4}$ ${\times}10^{-4} 1×1$${\times}10^{-4}$ $5{\times}10^{-5}$
Table 10: Optimal learning rates – lower parameter budget / non‑I.I.D. clients.
@@
Method ViT‑B‑16/CIFAR‑100 ViT‑B‑16/SVHN T5‑Base/20 Newsgroups T5‑Base/MRQA
FedIT $5{\times}10^{-3}$ ${\times}10^{-3} 1×1$${\times}10^{-3}$ $1{\times}10^{-3}$
FedEx‑LoRA ${\times}10^{-3} 1×1$${\times}10^{-3}$ ${\times}10^{-3} 1×1$${\times}10^{-3}$
FFA‑LoRA ${\times}10^{-2} 1×1$${\times}10^{-2}$ $1{\times}10^{-2}$ $5{\times}10^{-3}$
Fed‑SB $5{\times}10^{-4}$ $5{\times}10^{-3}$ $1{\times}10^{-3}$ $5{\times}10^{-4}$
@@
Table 11: Optimal learning rates – higher parameter budget / non‑I.I.D. clients.
@@
Method ViT‑B‑16/CIFAR‑100 ViT‑B‑16/SVHN T5‑Base/20 Newsgroups T5‑Base/MRQA
FedIT $5{\times}10^{-3}$ ${\times}10^{-3} 1×1$${\times}10^{-3}$ $5{\times}10^{-4}$
FedEx‑LoRA $1{\times}10^{-3}$ $5{\times}10^{-4}$ $1{\times}10^{-3}$ $5{\times}10^{-4}$
FFA‑LoRA ${\times}10^{-2} 1×1$${\times}10^{-2}$ $1{\times}10^{-2}$ $5{\times}10^{-3}$
Fed‑SB $5{\times}10^{-4}$ ${\times}10^{-3} 1×1$${\times}10^{-3}$ $5{\times}10^{-4}$
RAVAN $5{\times}10^{-4}$ $5{\times}10^{-4}$ ${\times}10^{-4} 1×1$${\times}10^{-4}$
@@
Figure 5: Fraction of clients assigned to each trainable parameter budget in each distribution. Skewed‑left is omitted because it is never used in our experiments.
To emulate computational heterogeneity in our FL setup, we vary the number of trainable parameters at each client. Let $N_{\max}$ denote the largest number of trainable parameters that any client can afford. Each client $c$ is constrained to a trainable parameter budget $N_{c}\in\{\frac{1}{4},\frac{1}{2},\frac{3}{4},1\}N_{\max}$. This value is held constant throughout the FL procedure to mirror fixed hardware limits. $N_{c}$ simply determines the trainable parameter budget per-client; all other hyperparameters are identical to the compute-homogeneous experiments. The bar plot in Figure 5 provides a visual summary of the client mix. Uniform serves as a neutral baseline where there is an equal proportion of clients at every trainable parameter budget; bell-shaped concentrates clients around medium ranks, reflecting the case where most devices have moderate capability; skewed‑right stresses the system by placing a large share of the population at the lowest rank, leaving only a small fraction of high‑capacity contributors. HetLoRA, FlexLoRA, and Ravan each accommodate clients with different trainable‑parameter budgets in distinct ways:
  • HetLoRA and FlexLoRA: The LoRA rank for each client $c$ is $r_{c}=\bigl\lfloor\left(N_{c}\,/\,N_{\max}\right)\cdot r_{\max}\bigr\rfloor$ where $r_{\max}$ is the maximum rank trained by any client. Since the number of trainable parameters scales linearly with the rank, this scaling ensures that every client keeps its update within the allotted budget $N_{c}$ while allowing higher‑capacity devices to contribute proportionally higher-rank updates.
@@
ViT‑B‑16 query,  value
T5‑Base SelfAttention.q,  SelfAttention.v
LLaMA3.2‑1B q_proj,  v_proj
For every model backbone, we insert LoRA adapters only in the self‑attention projection matrices. The exact parameters for which we apply LoRA are described in Table 12. All other parameters are frozen and do not have associated LoRA parameters.
@@ A.4  Additional Experiments
(a) CIFAR‑100 / lower parameter budget
(b) CIFAR‑100 / higher parameter budget
@@
Figure 6: Impact of computational heterogeneity on baselines and Ravan across four datasets. Each row groups two‑budget plots for the datasets listed in order.
@@
Method MNLI-$M$ MNLI-M QNLI QQP SST-2 RTE Average
@@
Method MNLI-$M$ MNLI-M QNLI QQP SST-2 RTE Average
@@
Method MNLI-$M$ MNLI-M QNLI QQP SST-2 RTE Average
@@
Method MNLI-$M$ MNLI-M QNLI QQP SST-2 RTE Average
@@
(a) CIFAR‑100 / 20 clients / lower parameter budget
(b) CIFAR‑100 / 50 clients / lower parameter budget
(c) CIFAR‑100 / 20 clients / higher parameter budget
(d) CIFAR‑100 / 50 clients / higher parameter budget
@@
Figure 7: Federated‑learning training curves for CIFAR‑100 and SVHN under different client counts and parameter‑budget settings.